Shift Summary, 10/11/99 Eve Shift

Arun Tripathi (arun@physics.ucla.edu)
Mon, 11 Oct 1999 22:15:31 -0700 (PDT)


Shift Summary for Shift, 10/11/99 Eve

On Shift: R. Tesarek/A. Tripathi

Hours of beam available at all intensities: 4hrs 30min
Hours of beam available at nominal intensity: ~4hrs 10min
Hours of E799 data (special runs not included): 3hrs 40min
Hours of special runs (laser scans, muon runs...): ~15 min

Accounting of lost time: TeV turnon, Initializing Beam_799 run after
the muon run, kfrending and resetting the
pipeline during the muon run, etc.

========================================================================

Accesses during shift: None
Runs Ended this Shift: 14820, 14821 - BEAM_MUON

========================================================================

Overview of shift: We finally get beam in ages! Overall smooth data
taking, no daq_dead even at 9.2E12. New CSQ6
constants installed from Peter Shanahan. Arun and
BobT look at random L3 events from run 14794 for
further insight into the L3 accept fraction problem.
Nothing obviously wrong is found in these events.
When I ran the filter code on events before and
after spill 48 (where problem started), I get
consistent results. BobT did not notice anything
funny in Kdisp.
The new alarm about L3 accept problems works, as
we found out during the muon run. Not a real problem
though. L3 accept fractions look great for BEAM_799
run 14822 so far!

========================================================================

Description of Shift:

run spill time comment
-----------------------------------------------------------------
17:40 Peter calls to say he is going to
install new DPMT constants.
14820 18:25 MCR threatens us with beam within the
hour. Intialize BEAM_MUON run 14820.
Install new CSQ6 constants upon Peter's
instructions.
14820 24 18:58 We get BEAM! Goes away after two spills.
43 Beam Back, ratty structure, intermittent.
19:15 TRD2 alarm comes on. ~128 uA current on
anode B. Goes away on its own quickly.
19:18 Sasha L. arrives and investigates the
TRD problem.
19:33 Sydney warns us about L3 accept fraction
problems. Not real though, the level
of the green histos for the muon runs
varies quite a bit across different CPU's.
But the ratios of green/red look OK.
14822 20:20 Started BEAM_799 run 14822.
Beam off target for spills 19-25. MCR
tuning the beam.
23:40 Beam intensity down to ~4E12. Comes back
to nominal after 4-5 spills.
00:00 Over to owl people.

========================================================================
Items to Follow up on Next Shift:

Load new T0's before starting a new run.